OpenStudy (anonymous):

Can someone help me simplify this rational expression? (Will post it below) THANK YOU!

OpenStudy (anonymous):



OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Ooo what a mess >.<

OpenStudy (anonymous):

I know :(

zepdrix (zepdrix):

\[\Large\rm \frac{\frac{3}{x}+\color{orangered}{\frac{1}{x^2+x}}}{\frac{1}{x+1}-\frac{1}{x-1}}\]So if you look at this orange part first... Hmm each term has something in common. We can factor the denominator, yes?

OpenStudy (anonymous):

Yes, the one in orange is (x+1)(x-1)

zepdrix (zepdrix):

Noooo, no no no :O

zepdrix (zepdrix):

\[\Large\rm (x+1)(x-1)=x^2-1\]

OpenStudy (anonymous):

hahahah, i suck at this. hellpppp :(

zepdrix (zepdrix):

They each have an x, yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So factor an x out :d \[\Large\rm x^2+x=x(x+1)\]

zepdrix (zepdrix):

\[\Large\rm \frac{\frac{3}{x}+\color{orangered}{\frac{1}{x(x+1)}}}{\frac{1}{x+1}-\frac{1}{x-1}}\]Yah? :d

OpenStudy (anonymous):

ohhhhh, right.. haha :P

zepdrix (zepdrix):

I'm not sure how you have been taught to do these..... but um.... I have learned .... in my years of doing math... that ... fractions are pure evil. So let's try something fancy. Here is what I wanna do...

zepdrix (zepdrix):

I wanna multiply the top and bottom by x(x+1)(x-1), just to get rid of every fraction.

zepdrix (zepdrix):

What'dyou think? Too fancy?

OpenStudy (anonymous):

That sounds like what my teacher was doing. So I think we can do it!

zepdrix (zepdrix):

x, (x+1) and (x-1) are the factors showing up in our denominators. \[\Large\rm \frac{\frac{3}{x}+\frac{1}{x(x+1)}}{\frac{1}{x+1}-\frac{1}{x-1}}\cdot\color{royalblue}{\left(\frac{x(x+1)(x-1)}{x(x+1)(x-1)}\right)}\]So by multiplying through by the LCM of our factors, we can get rid of these fractions.

zepdrix (zepdrix):

I wanna make sure you understand what we're doing. So... When we distribute this x(x+1)(x-1) to the first term:\[\Large\rm \frac{3}{x}x(x+1)(x-1)\]What happens to all that good stuff, where does it go? Numerator? Denominator of this fraction? What do you think?

OpenStudy (anonymous):

the x in the denominator and the x in the numerator cancel, and then the (x+1)(x-1) go on the numerator ? :) I hope haha

zepdrix (zepdrix):

Mmm ok good. So since we're left with nothing in the denominator, (Or a 1), we can simply write is as,\[\Large\rm 3(x+1)(x-1)\]ya? :d

OpenStudy (anonymous):

yes! I got that :)

zepdrix (zepdrix):

\[\large\rm \frac{3\cancel{x}(x+1)(x-1)}{\cancel{x}}=\frac{3(x+1)(x-1)}{1}=3(x+1)(x-1)\]

zepdrix (zepdrix):

Ok cool, do it with the othersssss +_+ Figure it outtttt You can do eeeet

OpenStudy (anonymous):

the other one on the top is \[1x(x+1)\]

OpenStudy (anonymous):

then on the bottom, \[1x(x-1) and 1(x-1)\]

zepdrix (zepdrix):

wait wait wait :U water you talking about? 1x(x+1)? I think you need to cancel out your x(x+1) from those factors, yes? leaving you with (x-1) for the second top term?

OpenStudy (anonymous):

ohhh, yes i see that now :P sorry haha

zepdrix (zepdrix):

\[\Large\rm \frac{3(x+1)(x-1)+(x-1)}{\frac{1}{x+1}-\frac{1}{x-1}}\cdot\left(\frac{1}{x(x+1)(x-1)}\right)\]And then what? You got some stuff in the bottom?

zepdrix (zepdrix):

The second term in the bottom looks messed up :U

OpenStudy (anonymous):

wait, can you explain why we are multiplying by (1/x(x+1)(x-1)) ?

zepdrix (zepdrix):

We already distributed the numerator, it's taken care of. So all that's left up there is a 1, we can't leave a 0 there.

zepdrix (zepdrix):

\[\Large\rm \frac{\frac{3}{x}+\frac{1}{x(x+1)}}{\frac{1}{x+1}-\frac{1}{x-1}}\cdot\left(\frac{1\cdot x(x+1)(x-1)}{x(x+1)(x-1)}\right)\]You can think of it like this if you like, with a 1 on top. After distributing the stuff to the numerator, and doing that fancy magic, we're left with: \[\Large\rm \frac{3(x+1)(x-1)+(x-1)}{\frac{1}{x+1}-\frac{1}{x-1}}\cdot\left(\frac{1}{x(x+1)(x-1)}\right)\]

OpenStudy (anonymous):

ohh okay.

OpenStudy (anonymous):

So how do I get rid of the fraction on the bottom? the same way as we did with the numerator?

zepdrix (zepdrix):

\[\Large\rm \frac{3(x+1)(x-1)+(x-1)}{\left(\frac{1}{x+1}-\frac{1}{x-1}\right)x(x+1)(x-1)}\cdot\left(\frac{1}{1}\right)\]Yah, distribute the stuff to each term. Your first fraction looked ok. You cancelled out the (x+1)'s.

zepdrix (zepdrix):

\[\Large\rm \frac{3(x+1)(x-1)+(x-1)}{x(x-1)-\color{green}{\frac{1}{x-1}~x(x+1)(x-1)}}\]How bout the second term? :d

OpenStudy (anonymous):

x(x+1)

zepdrix (zepdrix):

\[\Large\rm \frac{3(x+1)(x-1)+(x-1)}{x(x-1)-x(x+1)}\]Yay good job \c:/

zepdrix (zepdrix):

From there, you can uhhhh.. simplify if you want :d

OpenStudy (anonymous):

oh jeez.. haha. greeeaattt. Can i cancel things out like the (x+1) and (x-1) or should I distribute everything, combine like terms and then cancel things out?

zepdrix (zepdrix):

No canceling here :( Gotta distribute everything, then combine like-terms, then maybe if you're lucky, the leftover stuff will factor and you can try to factor maybe. as a matter of factor, i factor that you can factor those factors.. after you factor them of course. Ignore that last line ^ kinda lost my marbles there.

zepdrix (zepdrix):

then maybe if you're lucky, the leftover stuff will factor and you can try to `cancel`* maybe. Blah typo in there.

OpenStudy (anonymous):

okay, im getting \[3x ^{2}+x-4 \div -2\]

zepdrix (zepdrix):

ew ew ew D: you used the old school division operator! from kindergarden!! D:

OpenStudy (anonymous):

sorry, all of that is over -2 not 4 divided by -2. I have a feeling that is not right! GAH!

OpenStudy (anonymous):

I don't know how to use the line one! I'm not all fancy like you! :P

zepdrix (zepdrix):

Mmmmm yah I think that's right. I'm just eyeballing it, but it uhhhhhhh..... yah 3x^2+x-4 all over -2 wait wait wait wait.... .... !111!!!!1!one! isn't it -2x on the bottom?

OpenStudy (anonymous):

ahh, yes it is! Good catch! haha. Wait! my friend just messaged me and the teacher gave her the answer: (3x+4)(x-1) all over -2x so how do i get from 3x^2 +x -4 all over -2x to that ? D:

zepdrix (zepdrix):

So we need to factor the numerator :U

zepdrix (zepdrix):

friend to the rescue.

OpenStudy (anonymous):

ohh haha. it didn't send. I was trying to send "Just factor it?" okayyy.

OpenStudy (anonymous):

OH DANG. WE DID IT. hahah. Thank you so much!! You're wonderful!

zepdrix (zepdrix):

yay team \c:/ now get that dog out of my face -_- I'm allergic. No I'm not, but he looks like a racist. jk, im sure he's a sweetheart.

OpenStudy (anonymous):

awhh. Ben is the sweetest thing you could ever meet. He is such a lover. He just cuddles and kisses you all day.

zepdrix (zepdrix):

CATS RULE. <3 DOGS DROOL. there it is.

OpenStudy (anonymous):

I really am allergic to cats though D: